OpenStudy (anonymous):

A force of 10.0 newtons acts at an angle 20.0 degrees from vertical. What are the horizontal and vertical components of the force?

OpenStudy (anonymous):

A stone is dropped from a tower that is 100 meters high. The stone enters a well and continues its fall. If the stone takes 4 seconds to reach the surface of the water in the well, calculate how far the surface of the water is below ground level. Given: g = -9.8 meters/second2

OpenStudy (anonymous):

1. vertical component is 10 cos20 and horizontal component is 20 sin20

OpenStudy (anonymous):

2. calculate v by equation v=u+gt by putting u=0 and g=9.8 and t=4 after that calculate s by equation \[v^2 =2gh\] after that subtract 100 from h you will get final answer.